OpenStudy (anonymous):

A college student earned $5100 during summer vacation working as a waiter in a popular restaurant. The student invested part of the money at 9% and the rest at 8%. If the student received a total of $445 in interest at the end of the year, how much was invested at 9%?

OpenStudy (anonymous):

\[.09x+.08(2100-x)=445\]is the equation to solve

OpenStudy (anonymous):

oops \[.09x+.08(5100-x)=445\]

OpenStudy (anonymous):

or \[9x+8(5100-x)=44500\] if you do not feel like working with decimals